Algebra Find the missing measurement for each triangle.
A = 36 mi 2 , b = (blank) , h = 12 mi
A = (blank) , b = 12 mm, h = 7.5 mm

Answer by jim_thompson5910(35256)

I'll do the first one to get you started

A = (b*h)/2
36 = (b*12)/2
36*2 = 12b
72 = 12b
12b = 72
b = 72/12
b = 6

So b = 6 is the answer to the first question.
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------